# Automated radiotherapy treatment planning guided by GPT-4Vision

### v2

@@ 2.3 Dataset
This study utilizes a dataset comprising imaging and treatment plans for 11 prostate cancer patients and 6 head & neck cancer patients who underwent VMAT at the Stanford University Cancer Center. Available data for each patient includes CT scans, delineated anatomical structures, and clinically approved treatment plans obtained via Eclipse® (Varian Medical Systems, Palo Alto, CA).
For prostate cancer patients, all cases adhere to a clinical protocol designed to deliver a total dose of 70.2 Gy to the majority of the PTV. Concurrently, the protocol stipulates specific dose constraints for the OARs, which include the bladder, rectum, and both femoral heads. These dose limits typically involve upper caps on the mean OAR doses and DVH constraints, such as restricting no more than 34% of the rectum to receive in excess of 31 Gy.
Treatment plans for headand neck cancer patients follow a protocol that prescribes a dose of 70 Gy to the primary PTV. Additionally, these patients have two PTV volumes receiving doses of 56 Gy and 52 Gy, respectively. Key OARs for this group are the brainstem, spinal cord, oral cavity, larynx, and both parotid glands. The dose constraints for these OARs include upper limits on the mean dose received, with several organs also subjected to maximum dose limits.
@@ 2.4 Technical details and evaluation
We develop an in-house two-loop optimization system, where GPT-RadPlan performs outer loop optimization of the optimizer weights, and the inner loop fluence map optimization is based on the open-source TPS matRad Wieser et al. (2017). All plans utilize a CT and dose voxel grid with $2.5\times 2.5\times 2.5\text{mm}^{3}$ resolution. MatRad VMAT plans consist of co-planar arcs and are based on the SmartArc planning algorithm MacFarlane et al. (2020); Christiansen et al. (2018), using the interior-point optimization (IPOPT) package for the inverse optimization step.
Throughout our experiments, we compare the plans generated by GPT-RadPlan to their clinical counterpart, using multiple dose metrics and indexes that are frequently used in the clinic. These include the $D_{95}$, the homogeneity index (HI) and the conformity index (CI) and for the targets, and the $D_{5}$, $D_{50}$, $V_{15}$ and $V_{30}$ for the OARs.
@@
$HI=\frac{D_{5}-D_{95}}{d_{t}}\times 100,$ (2)
where lower values close to 0 indicate ideal dose homogeneity. Alternatively, the CI measures both the conformation of doses on target and the volume of surrounding tissue that is covered by the reference dose. While multiple methods have been proposed to calculate this CI, we opt for the following definition
@@
where $TV_{95,PTV}$ and $TV_{95,Body}$ denotes the total volumes receiving at least 95% of the prescribed dose for the PTV and the whole body, respectively, and $TV$ is the total body volume. In the ideal scenario, the volume receiving the prescribed dose is confined to just the PTV, and the CI tends to 1 as the numerator and denominator become equal.
@@ 3.1 Planning performance
To assess GPT-RadPlan capabilities in mimicking clinical human planners, we first perform a quantitative comparison between GPT-RadPlan generated plans and clinical plans based on the aforementioned metrics, including both the CI and HI, the $D_{95}$, mean and minimum doses for the PTV targets, and the mean dose, $D_{5}$, $D_{50}$, $V_{15}$ and $V_{30}$ for the OARs. We also compare the types of plans qualitatively based on the final DVH and dose distributions.
Table 1 contains the average PTV metrics across all plans for both disease sites. For the prostate cases, GPT-RadPlan’s results outperform the clinical plans across all metrics, achieving a higher minimum dose, a $D_{95}$ that exactly matches the prescribed dose, and significantly better dose homogeneity and conformity. Similarly, GPT-RadPlan demonstrated enhanced performance over the clinical plans in the head & neck cases, showing improved metrics across all categories, albeit with a slight lag in target homogeneity for the high-dose target (PTV70). Additionally, Table 2 displays the organ-sparing metrics for the main OARs. As in the PTV case, GPT-RadPlan results in treatment plans that deliver less dose to OARs, as shown by the lower metric values across the board.
Table 1: PTV dose metrics. Several dose metrics of the PTV target are displayed for all the clinical and GPT-RadPlan plans, including the mean and minimum doses, as well as the $D_{95}$, the HI and the CI. For all the metrics, we include the average values across plans and the standard deviation in brackets. Values in bold represent the best for each PTV target.
@@
PTV70.2 Goal 70.20 $\approx 70.20$ 70.20 0 1.00
Clinical 72.35 (0.63) 61.16 (2.39) 70.05 (0.15) 5.43 (1.04) 0.883 (0.021)
GPT-RadPlan 70.69 (0.09) 63.69 (2.33) 70.21 (0.01) 1.96 (0.26) 0.921 (0.025)
@@
PTV70 Goal 72.00 $\approx 70.00$ 70.00 0 1.00
Clinical 71.97 (0.23) 60.75 (3.38) 69.69 (0.08) 5.40 (0.51) 0.847 (0.018)
GPT-RadPlan 72.70 (0.26) 67.78 (1.06) 70.01 (0.01) 6.88 (0.76) 0.963 (0.004)
PTV56 Clinical 62.05 (1.32) 43.50 (8.41) 54.82 (2.96) 26.53 (6.91)
GPT-RadPlan 59.34 (0.56) 49.46 (2.32) 56.07 (0.24) 14.92 (1.21) -
PTV52 Clinical 54.11 (0.22) 44.49 (2.24) 52.23 (0.55) 6.83 (1.01)
GPT-RadPlan 53.92 (0.03) 46.43 (2.40) 52.74 (0.33) 3.91 (0.97) -
Table 2: OAR dose metrics. We show multiple dose metrics capturing OAR sparing in prostate and head & neck plans. These include the mean dose, the $D_{5}$ and $D_{50}$ (maximum common dose that 5% or 50% of the volume receives), and the $V_{15}$ and $V_{30}$ (representing which percentage of the volume receives 15 and 30 Gy, respectively). For all the metrics, we include the average values across plans and the standard deviation in brackets. In all cases, lower values are preferred.
@@
Bladder Clinical 20.42 (8.87) 63.60 (11.67) 13.18 (11.04) 42.40 (17.21) 27.56 (14.20)
GPT-RadPlan 19.28 (7.90) 64.14 (11.04) 11.20 (8.56) 40.11 (14.84) 26.07 (12.53)
Rectum Clinical 26.84 (6.43) 66.92 (6.91) 22.79 (6.64) 68.40 (15.46) 36.03 (10.64)
GPT-RadPlan 22.56 (4.46) 62.57 (10.97) 16.54 (4.46) 53.91 (12.68) 26.98 (9.44)
@@
Brainstem Clinical 12.25 (3.46) 31.36 (2.30) 8.60 (5.01) 32.90 (13.70) 9.74 (7.37)
GPT-RadPlan 11.49 (4.76) 27.92 (3.24) 9.10 (6.70) 35.44 (16.31) 0.00 (0.00)
Larynx Clinical 46.30 (12.02) 72.30 (0.60) 44.28 (16.52) 96.27 (3.44) 69.04 (20.11)
GPT-RadPlan 40.74 (14.49) 73.16 (0.36) 35.08 (22.22) 76.48 (6.69) 55.58 (25.50)
Oral Cavity Clinical 35.55 (1.73) 70.43 (3.31) 40.55 (18.14) 93.68 (4.46) 50.57 (3.84)
GPT-RadPlan 31.26 (0.71) 69.89 (4.93) 24.35 (2.55) 75.60 (9.74) 38.29 (0.49)
Parotid Left Clinical 36.36 (14.22) 61.16 (12.28) 37.35 (21.66) 74.80 (20.07) 54.53 (22.41)
GPT-RadPlan 30.31 (15.05) 58.28 (15.80) 31.74 (22.77) 56.81 (22.79) 39.27 (25.08)
Parotid Right Clinical 34.67 (9.31) 60.88 (10.68) 33.09 (12.09) 81.64 (19.21) 54.86 (19.70)
GPT-RadPlan 27.42 (3.65) 60.39 (10.76) 21.64 (5.46) 61.06 (11.23) 36.74 (8.55)
Spinal Cord Clinical 17.36 (6.74) 33.25 (5.66) 17.13 (13.53) 54.81 (12.25) 22.60 (27.89)
GPT-RadPlan 15.67 (3.08) 30.23 (0.19) 14.26 (11.72) 54.31 (11.23) 0.00 (0.00)
@@
Figure 2 further confirms GPT-RadPlan’s superiority in sparing OARs. In the prostate cases depicted in Figure 3(a), GPT-RadPlan achieves a sharper dose fall-off within the PTV (black line), resulting in more homogeneous delivery. Additionally, GPT-RadPlan generates plans that considerably spare the rectum compared to their clinical counterparts, while maintaining comparable doses to the femoral heads and slightly decreasing average bladder doses. In the head & neck cases illustrated in Figure 3(b), the clinical plans generally result in lower doses to all OARs, particularly for the larynx, oral cavity, and parotid glands, which often partially overlap with the target volumes.
Due to the overlapping standard deviations noted in Table 2 and the wide shaded regions in Figure 2, GPT-RadPlan’s superiority over clinical plans in sparing OARs remains uncertain. The large standard deviations occur because the tumor targets in each patient are located in closer proximity to different OARs. For instance, two head & neck patients might each have a larger PTV overlapping either the left or right parotid gland, leading to a substantially higher dose received by the overlapping gland. We demonstrate this and prove GPT-RadPlan’s superiority in Figure 3, where the difference in dose metrics between GPT-RadPlan and clinical plans is calculated individually for each patient in the cohort. The data consistently show that GPT-RadPlan’s metrics are lower, as the boxes generally sit above the red line, confirming that GPT-RadPlan is capable of achieving plans that deliver less dose on the same patient. Overall, GPT-RadPlan plans outperforms clinical plans in 82% of the prostate patients, achieving an average 15% reduction in OAR metrics. Likewise, GPT-RadPlan provides better plans in 75% of the head & neck cases, reducing the mean dose by a 10-15% for most of the OARs.
@@
Figure 3: Case-by-case differences between GPT-RadPlan and clinical plans. For every relevant OAR, each box displays the relative difference between clinical and GPT-RadPlan plans of all the patients in the prostate or head & neck cohorts. Each panel shows the relative difference between different OAR metrics, including (left) the mean dose, (center) the $D_{50}$ and (right) the $V_{15}$. Positive values above the red line indicate that GPT-RadPlan plans result in lower metric values, which is preferred.
To investigate how GPT-RadPlan acts over the multiple iterations, Figure 4 compares the dose distributions and DVHs for a prostate and head & neck patient at different iteration steps. Figure 4(a) depicts such “planning trajectors” for one of the prostate patients, where GPT-RadPlan first ensures homogeneous PTV coverage (uniform red color) and subsequently reduces the dose delivered to the rectum and bladder (blue and red contours, respectively). Likewise, Figure 4(b) demonstrates the same behavior for a head & neck patient, where GPT-RadPlan first ensures target coverage (black, purple, and magenta contours), and then progressively minimizes the dose delivered to the oral cavity and larynx (grey and green contours), as well as the overall body dose.

### v3

@@ 2.3 Dataset
This study utilizes a dataset comprising imaging and treatment plans for 17 prostate cancer patients and 13 head & neck cancer patients who underwent VMAT. Available data for each patient includes CT scans, delineated anatomical structures, and clinically approved treatment plans created via Eclipse® (version 15.6, Varian Medical Systems, Palo Alto, CA).
For the prostate cancer patients, all cases adhere to a clinical protocol designed to deliver the mean dose of 70.2 Gy to the PTV. Concurrently, the protocol stipulates specific dose constraints for the OARs, which include the bladder, rectum, and both femoral heads. These dose limits typically involve constraints on the mean OAR doses and DVH constraints, such as restricting no more than 34% of the rectum to receive in excess of 31 Gy.
Treatment plans for the head & neck cancer patients follow a protocol that prescribes a dose of 70 Gy to the primary PTV (PTV70). Additionally, two nodal PTVs — PTV56 and PTV52 — receive prescribed doses of 56 Gy and 52 Gy, respectively. All these doses are prescribed to at least 95% of the target volume. Key OARs for this group are the brainstem, spinal cord, oral cavity, larynx, and both parotid glands. The dose constraints for these OARs include upper limits on the mean dose received, with several organs also subjected to maximum dose limits.
@@ 2.4 Technical details and evaluation
We develop an in-house two-loop optimization system, where GPT-RadPlan performs outer loop optimization of the optimizer weights, and the inner loop fluence map optimization is based on the open-source treatment planning system (TPS) matRad [31]. All plans utilize a CT, and the dose grid had a voxel size of $2.5\times 2.5\times 2.5\text{mm}^{3}$ resolution. MatRad VMAT plans consist of 1 to 3 full co-planar arcs and are based on the SmartArc planning algorithm [32, 33], using the interior-point optimization (IPOPT) package for the inverse optimization step.
Throughout our experiments, we compare the plans generated by GPT-RadPlan to plans optimized by multiple baseline methods and human-designed clinically approved plans. These include the plans generated by certified physicists from our institution with multiple years of experience in developing treatment plans across various cancer types, including prostate and head & neck cases. We also include plans obtained as a result of running an automated planning algorithm [34], referred to as Autoplan in the remainder of the manuscript, which is a knowledge-guided iterative algorithm based on a dose-anatomy guided voxelization scheme. The Autoplan script first sets up all the treatment planning parameters automatically based on the prescription and disease sites. The prediction model based on the dosimetric metric from the simple treatment plans (PTV-only plans) is then used to predict and initialize the personalized optimization constraints automatically for a new patient with the script. The script includes a plan evaluator for quantitative plan assessment and the algorithm to automatically update treatment plan parameters in each iteration. These perform experience-based adjustments, first checking the dose/volume with the preset constraints and subsequently modifying the weights and dose constraints based on difference between the results of the current iteration and the clinical requirements [34].
The third baseline, referred to as Bayesian Optimization (BO), where we adopted the plan quality metrics (PQM) mentioned in [35] as the objective in the outer loop and used Optuna [36], a commonly used package designed for Bayesian optimization, to implement Bayesian optimization for optimizing parameters. Since the original paper only presented this BO method and provided the model configuration for prostate cases, we only apply BO to our prostate cohort.
To compare plans, we use multiple dose metrics and indexes that are frequently used in the clinic. These include the the mean dose, $D_{95}$, homogeneity index (HI) and conformity index (CI) and for the targets, and the $D_{5}$, $D_{50}$, $V_{15}$ and $V_{30}$ for the OARs.
@@
$HI=\frac{D_{5}-D_{95}}{d_{p}}\times 100,$ (2)
where lower values close to 0 indicate ideal dose homogeneity. The CI measures both the conformation of doses on target and the volume of surrounding tissue that is covered by the reference dose. While multiple methods have been proposed to calculate this CI, we opt for the following definition
@@
where $TV_{95,PTV}$ and $TV_{95,Body}$ denotes the total volumes receiving at least 95% of the prescribed dose for the PTV and the whole body, respectively, and $TV$ is the total target volume. In the ideal scenario, the volume receiving the prescribed dose is confined to just the PTV, and the CI tends to 1 as the numerator and denominator become equal.
@@ 3.1 Planning performance
To assess GPT-RadPlan capabilities in mimicking clinical human planners, we first perform a quantitative comparison between GPT-RadPlan generated plans and clinical plans based on the aforementioned metrics. We also compare the types of plans based on the final DVH and dose distributions.
Table 1 contains the average PTV metrics across all plans for both disease sites. For the prostate cases, GPT-RadPlan outperforms all baselines including clinical human planners across all metrics, achieving a higher minimum dose, a $D_{95}$ that exactly matches the prescribed dose, and significantly lower compared dose homogeneity and conformity. In the head & neck cases, GPT-RadPlan matches or outperforms the baselines, showing higher minimum doses and better conformity. Additionally, Table 2 displays the organ-sparing metrics for the main OARs. For both the prostate and head & neck cases, GPT-RadPlan produces treatment plans that generally deliver lower doses to OARs compared to the baselines, as reflected by improved values in most evaluation metrics.
Table 1: PTV dose metrics. Several dose metrics of the PTV target are displayed for all the GPT-RadPlan plans and baselines, including the mean and minimum doses, as well as the $D_{95}$, the HI and the CI. For all the metrics, we include the average values across plans and the standard deviation in brackets. Values in bold represent the statistically best result for each PTV target, when the average minus the standard deviation is higher than the average plus the standard deviation of the second best value.
@@
PTV70.2 Clinical 72.47 (0.64) 61.34 (2.21) 70.04 (0.18) 5.43 (1.04) 0.883 (0.020)
Autoplan 72.78 (0.52) 61.83 (1.24) 70.03 (0.28) 6.01 (0.60) 0.870 (0.011)
BO 73.04 (1.65) 60.66 (2.28) 69.76 (0.46) 5.05 (0.83) 0.875 (0.016)
GPT-RadPlan 70.94 (0.65) 63.82 (2.19) 70.21 (0.06) 2.43 (1.11) 0.912 (0.041)
@@
PTV70 Clinical 71.85 (0.26) 60.57 (3.72) 69.69 (0.14) 5.05 (0.58) 0.881 (0.048)
Autoplan 72.05 (0.27) 62.12 (2.73) 69.82 (0.17) 5.01 (0.47) 0.871 (0.038)
GPT-RadPlan 72.73 (0.30) 62.73 (1.10) 70.00 (0.01) 7.12 (0.71) 0.948 (0.005)
PTV56 Clinical 65.47 (3.53) 41.40 (9.71) 56.43 (2.09) - -
Autoplan 67.58 (2.71) 49.07 (10.9) 57.45 (0.60) - -
GPT-RadPlan 64.66 (4.51) 49.62 (2.18) 57.31 (1.39) - -
PTV52 Clinical 54.58 (0.80) 44.57 (2.01) 52.14 (0.53) - -
Autoplan 55.44 (0.44) 44.08 (1.73) 51.89 (0.46) - -
GPT-RadPlan 54.26 (0.45) 45.04 (3.09) 52.42 (0.24) - -
Table 2: OAR dose metrics. Multiple dose metrics capturing OAR sparing in prostate and head & neck plans are presented. These include the mean dose, the $D_{5}$ and $D_{50}$ (maximum common dose that 5% or 50% of the volume receives), and the $V_{15}$ and $V_{30}$ (representing which percentage of the volume receives 15 and 30 Gy, respectively). For all the metrics, we include the average values across plans and the standard deviation in brackets. In all cases, lower values are preferred.
@@
Bladder Clinical 24.48 (9.30) 62.14 (10.79) 18.87 (12.12) 60.85 (29.52) 34.45 (15.28)
Autoplan 31.91 (1.11) 59.48 (7.71) 29.29 (0.87) 94.67 (3.92) 47.08 (2.72)
BO 23.88 (6.45) 64.26 (10.00) 20.04 (5.50) 54.13 (21.17) 35.18 (12.11)
GPT-RadPlan 17.02 (8.11) 60.08 (14.39) 9.18 (8.63) 35.60 (17.07) 22.62 (13.12)
Rectum Clinical 26.45 (5.64) 63.64 (8.82) 22.20 (5.71) 71.91 (14.92) 32.87 (10.54)
Autoplan 25.73 (3.19) 57.63 (8.76) 21.13 (2.26) 78.35 (10.45) 27.07 (6.66)
BO 22.39 (5.55) 67.81 (6.44) 22.13 (8.42) 75.32 (13.17) 28.14 (9.13)
GPT-RadPlan 21.80 (6.60) 58.88 (13.81) 17.09 (7.99) 50.44 (20.03) 27.65 (16.43)
@@
Brainstem Clinical 12.17 (4.29) 30.95 (4.78) 8.34 (5.75) 33.36 (14.10) 10.03 (11.03)
Autoplan 9.85 (3.41) 27.47 (5.84) 6.24 (3.42) 25.22 (9.81) 5.67 (5.86)
GPT-RadPlan 10.95 (4.89) 26.91 (4.74) 8.09 (7.51) 32.98 (16.07) 0.03 (0.11)
Larynx Clinical 42.80 (11.03) 69.95 (4.30) 39.85 (14.87) 97.44 (3.24) 64.06 (18.18)
Autoplan 41.89 (9.95) 69.08 (4.79) 37.13 (13.30) 99.94 (0.17) 60.92 (15.95)
GPT-RadPlan 34.15 (14.33) 70.68 (5.75) 26.96 (21.64) 64.12 (26.95) 42.46 (25.01)
Oral Cavity Clinical 34.86 (4.15) 68.07 (8.83) 33.62 (12.29) 92.96 (5.28) 50.00 (11.01)
Autoplan 35.61 (4.92) 68.08 (9.77) 30.96 (4.59) 97.39 (3.65) 51.06 (13.87)
GPT-RadPlan 25.41 (8.51) 66.38 (15.18) 18.24 (8.33) 54.32 (25.66) 27.89 (13.78)
Parotid Left Clinical 35.10 (13.93) 63.18 (10.85) 33.50 (21.21) 71.64 (20.97) 52.80 (24.12)
Autoplan 31.70 (11.77) 63.34 (10.52) 30.12 (20.03) 61.94 (16.5) 44.58 (19.62)
GPT-RadPlan 24.26 (12.65) 57.76 (16.06) 19.73 (18.61) 42.99 (20.67) 28.49 (21.04)
Parotid Right Clinical 34.32 (10.67) 61.69 (11.30) 32.22 (13.83) 79.00 (21.05) 52.65 (20.65)
Autoplan 29.93 (9.81) 61.17 (12.90) 27.51 (14.09) 62.63 (13.03) 41.85 (17.51)
GPT-RadPlan 21.66 (7.79) 58.49 (15.56) 13.66 (7.45) 41.84 (19.71) 25.60 (13.81)
Spinal Cord Clinical 17.36 (6.32) 33.88 (5.15) 18.84 (13.24) 55.32 (13.69) 24.83 (25.69)
Autoplan 14.91 (3.04) 32.07 (1.62) 13.74 (9.21) 53.00 (14.41) 16.00 (7.62)
GPT-RadPlan 16.27 (3.96) 30.14 (0.24) 17.76 (12.52) 55.78 (14.19) 1.34 (3.25)
@@
Figure 3: Average DVH comparison. Visual comparison of the average DVH lines across all prostate (left) and lung (right) patients. Solid lines represent GPT-RadPlan plans, while dashed lines indicate clinical plans.
Figures 2 and 3 further confirms GPT-RadPlan’s superiority in sparing OARs. In the prostate cases depicted in Figure 2(a), GPT-RadPlan achieves a sharper dose fall-off within the PTV (black line), resulting in more homogeneous delivery. Additionally, GPT-RadPlan generates plans that considerably spare the rectum compared to their clinical counterparts, while maintaining comparable doses to the femoral heads and slightly decreasing average bladder doses. In the head & neck cases illustrated in Figure 2(b), the clinical plans generally result in lower doses to all OARs, particularly for the larynx, oral cavity, and parotid glands, which often partially overlap with the target volumes. While the shaded areas in the clinical plans may indicate dose spillage in some head & neck plans, the physician’s intent only specifies a lower dose limit for PTV56 and PTV52, with no upper limit. While the human planners primarily focused on enforcing this lower limit, GPT-RadPlan employs an optimization objective that penalizes both positive and negative deviations from the 56 Gy target, which likely explains the differences observed in the DVH distribution.
To statistically demonstrate the superiority of our method over clinical plans in reducing OAR mean doses, we performed a Wilcoxon signed-rank test, a non-parametric test for paired samples which is commonly used to prove whether paired measurements are different without assuming normal distribution of the data. This test evaluates whether the median of the mean doses extracted from the different plans obtained from GPT-RadPlan are significantly lower than those of the clinical plans. Specifically, a one-tailed test was conducted with the null hypothesis stating that GPT-RadPlan method does not achieve a lower mean dose compared to clinical plans. The test provides p-values, and the null hypothesis is rejected when the p-value is less than the 0.05 significance level, which indicates that GPT-RadPlan achieves statistically significant improvements in mean dose reduction. After running the Wilcoxon test, we obtain a p-value of 0.003 and 0.001 for the rectum and bladder, demonstrating that GPT-RadPlan delivers less dose to these organs. Similar results are obtained for the head & neck cases, with p-values of 0.06, 0.006, 0.0001, 0.0001, 0.0003 and 0.729 for the brainstem, larynx, oral cavity, left and right parotid and spinal cord, respectively. These values further confirm that GPT-RadPlan can reduce the dose to most OARs except for the brainstem and spinal cord.
Due to the overlapping standard deviations noted in Table 2 and the wide shaded regions in Figure 2, GPT-RadPlan’s superiority over clinical plans in sparing OARs remains uncertain. The large standard deviations occur because the tumor targets in each patient are located in closer proximity to different OARs. For instance, two head & neck patients each have a larger PTV overlapping either the left or right parotid gland, leading to a substantially higher dose received by the overlapping gland. We demonstrate this and prove GPT-RadPlan’s superiority in Figure 4, where the difference in dose metrics between GPT-RadPlan and clinical plans is calculated individually for each patient in the prostate (Figure 4(a)) and head & neck (Figure 4(b))cohorts. The data show that GPT-RadPlan’s metrics are lower, as the boxes generally sit above the red line, confirming that GPT-RadPlan is capable of achieving plans that deliver less dose on the same patient. In general, the GPT-RadPlan plans reduce OAR doses while mantaining coverage in 82% of prostate patients, achieving an average reduction 5 Gy decrease in mean dose and 15% overall in OAR metrics as shown in 4(a). Likewise, GPT-RadPlan improves OAR sparing in 75% of the head & neck cases, reducing the mean dose by 4-6 Gy or 10-15% for most of the OARs.
@@
Figure 4: Differences between GPT-RadPlan and clinical plans. For every relevant OAR, each box displays the relative difference between clinical and GPT-RadPlan plans of all the patients in the prostate or head & neck cohorts. Each panel shows the relative difference between different OAR metrics, including (left) the mean dose, (center) the $D_{50}$ and (right) the $V_{15}$. Positive values above the red line indicate that GPT-RadPlan plans result in lower metric values, which is preferred.
Regarding the non-clinical baselines, GPT-RadPlan also yields better plans than the BO and Autoplan, as demonstrated by the lower OAR metrics in Table 2. In particular, Autoplan appears to perform better than BO, similar to clinical planners and closer to GPT-RadPlan. For this reason, we have included additional results in Appendix B.2. As shown in 8, although Autoplan often yields plans that are slightly worse than GPT-RadPlan’s results, often with higher OAR doses except for select organs such as the femoral heads or the brainstem. The relative differences in Figure 9 in the Appendix, showing overall values that fall above the red line, further confirm GPT-RadPlan’s superiority.
To investigate how GPT-RadPlan acts over the multiple iterations, Figure 5 compares the dose distributions and DVHs for a prostate and head & neck patient at different iteration steps. Figure 5 depicts such planning trajectories for one of the prostate patients, where GPT-RadPlan first ensures homogeneous PTV coverage (uniform red color) and subsequently reduces the dose delivered to the rectum and bladder (blue and red contours, respectively). Likewise, Figure 5 demonstrates the same behavior for a head & neck patient, where GPT-RadPlan first ensures target coverage (black, purple, and magenta contours), and then progressively minimizes the dose delivered to the oral cavity and larynx (grey and green contours), as well as the overall body dose. Most importantly, GPT-RadPlan takes 3-6 iterations to obtain the final plan (around 2-3 hours without human supervision needed), which is comparable to human planners and lower than the 7-10 iterations for Autoplan and more than 50 iterations for the BO baseline.